Subscribe to Standard Changes for IS Services and Systems feed
Announcements for OSU's Network and Computer Services
Updated: 1 hour 46 sec ago

A016 Data Center (May 16th 00:01 am)

Wed, 05/13/2015 - 11:54am
Primary Contact: Joel Burks, We will be upgrading the Nexus devices in a016. We have sent a notice to DCA-all with the details. Will be about a 20 minute outage.

DNS move to new network core

Mon, 05/11/2015 - 1:31pm
Primary Contact: Mike Akey, When: May 19th, 2015, 6AM We will be moving one of our name servers, ns1, to the new network core by way of moving it’s subnet to the new core. This will result in a brief service blip while the transition happens. No client impact is expected. This move should [...]

Firewall Maintenance

Mon, 05/11/2015 - 12:26pm
Primary Contact: Joel Burks, a redundant link went down a few months ago. We have moved off the bad fiber. Tomorrow at 6:15 am I’ll bring the link back up.

LS Networks planned maintenance on 5/21/2015 00:01 to 03:00

Mon, 05/11/2015 - 12:26pm
Primary Contact: Joel Burks, Expected Outage/Downtime: Expected outage up to 15 minutes. LSN will be performing maintenance on May 21st, 2015 00:01 to 03:00 PST in Corvallis, OR. LSN will be performing network groom to core equipment. Our Network Operation Center technicians will closely monitor the work and will do everything possible to minimize [...]

Upgrading Snell building switch Saturday 10pm-12am

Fri, 05/08/2015 - 10:23am
Primary Contact: Tim Bateman, Upgrading the building switch for Snell starting at 10pm Saturday, finishing by midnight.

Milne Data Center Maintenance April 14th 6 to 7 AM

Wed, 04/08/2015 - 3:29pm
Primary Contact: Joel Burks, We will be upgrading the Milne Nexus switches. A notice was sent to dca-all with details.

Banner web server patching

Tue, 03/24/2015 - 1:39pm
Primary Contact: Jill Swenson, We will be applying standard update patches to a Banner middle tier web server on March 24 from 6-7PM. A redundant web server will serve user requests during this maintenance. Some users may see a “FRM-92150: Fatal error: web client version is too new” message when starting Banner INB after [...]

Milne Data Center Maintenance March 24th 6 to 7 AM

Thu, 03/19/2015 - 10:09am
Primary Contact: Joel Burks, On March 24th at 6 AM the Network Operations team will be doing maintenance on two redundant switches in the Milne Data Center. The goal of this maintenance is to upgrade the firmware and implement best practices settings. This maintenance will be performed with Cisco TAC online. During the maintenance, [...]

FormFusion Maintenance

Mon, 03/09/2015 - 8:45am
Primary Contact: Jill Swenson, FormFusion will be unavailable from 5-6PM on Tuesday March 10 due to scheduled maintenance. We will be making configuration changes. FormFusion users have been informed of this maintenace through the FormFusion users mail list.

Banner Table Change

Wed, 02/25/2015 - 12:26pm
Primary Contact: Jill Swenson, On March 3 2015, between 9:00AM and 1:00PM, we will be redirecting a number of processes to point to a table which has been restructured (old table: SERSSGP, new table SYRSSGP). This change has been tested in DEVL. Updates include replacing one Banner INB form (SEAASGN will be replaced with [...]

Milne Data Center Maintenance March 3rd 6 to 7 AM

Tue, 02/24/2015 - 11:22am
Primary Contact: Joel Burks, We will be upgrading the Nexus 6000s in the Milne Data Center. Here is a list of the devices connected to the network gear. SIG Eth1/4 c7k-mcc Bay 1 Port Eth1/6 #datadomain link1 Eth1/23 #mcc-b5-sw39.sig l Eth1/25 Po308 mlag1a-nic1. Eth1/26 Po308 mlag1b-nic1. Po #datadomain Po308 #mlag-domain1 Po309 #mcc-b5-sw39.sig o Eth105/1/10 [...]

Data Center Maintenance in B210 (Feb 24th 6 am to 7 am)

Wed, 02/18/2015 - 12:00pm
Primary Contact: Joel Burks, B210 Data Center will be undergoing maintenance between 6 and 7 am.

Data Center Maintenance in B210 (Feb 17th 6 am to 7 am)

Wed, 02/11/2015 - 3:26pm
Primary Contact: Joel Burks, The data center in Kerr B210 will be undergoing maintenance on Feb 17th 6 to 7 am.

Data Center Maintenance in B210, A016 Kerr, and Milne

Tue, 02/03/2015 - 7:02am
Primary Contact: Joel Burks, Changing MTU size on 4 links, turning off arp sync on router, and implementing new management vlan. No outage expected.